# PrivATE: Differentially Private Confidence Intervals for Average Treatment Effects

### v1

@@ Abstract
The average treatment effect (ATE) is widely used to evaluate the effectiveness of drugs and other medical interventions. In safety-critical applications like medicine, reliable inferences about the ATE typically require valid uncertainty quantification, such as through confidence intervals (CIs). However, estimating treatment effects in these settings often involves sensitive data that must be kept private. In this work, we present PrivATE, a novel machine learning framework for computing CIs for the ATE under differential privacy. Specifically, we focus on deriving valid privacy-preserving CIs for the ATE from observational data. Our PrivATE framework consists of three steps: (i) estimating a differentially private ATE through output perturbation; (ii) estimating the differentially private variance through a truncated output perturbation mechanism; and (iii) constructing the CIs while accounting for the uncertainty from both the estimation and privatization steps. Our PrivATE framework is model agnostic, doubly robust, and ensures valid CIs. We demonstrate the effectiveness of our framework using synthetic and real-world medical datasets. To the best of our knowledge, we are the first to derive a general, doubly robust framework for valid CIs of the ATE under ($\varepsilon,\delta$)-differential privacy.
@@ 1 Introduction
Estimating the average treatment effect (ATE) from observational data is highly relevant for evaluating the effectiveness of drugs and other medical interventions [e.g., 7, 13, 25]. To ensure the reliability of ATE estimates, one often needs to “move beyond the mean” by accounting for the estimation uncertainty [31, 43]. This is typically achieved through confidence intervals (CIs), which provide a range within which the true effect is likely to lie [49]. Medical studies frequently report 95% CIs to show the precision of the estimated treatment effects. For example, the clinical trial for the Moderna COVID-19 vaccine specified that the effectiveness must meet a predefined threshold based on the 95% CI, and, ultimately, the trial reported a 95% CI of 89.3% to 96.8% [6].
Estimating ATEs in medicine typically involves sensitive patient data [12], that must be kept private. Many regulations have been introduced to enforce data privacy, such as the US Health Insurance Portability and Accountability Act (HIPAA) and the EU General Data Protection Regulation (GDPR). One approach to analyse sensitive data is to use privacy mechanisms such as differential privacy (DP), enabling population-level inference while protecting individual information [20]. Due to strong theoretical guarantees, especially ($\varepsilon,\delta$)-DP has received increasing attention [e.g., 1, 10].
In this work, we aim to compute differentially private CIs for the ATE. However, this is non-trivial: constructing CIs solely based on the private ATE estimate does not suffice to ensure DP of the entire CI, as estimating the variance required for constructing CIs involves additional queries to the underlying data. Thus, it is necessary to develop privacy-preserving mechanisms not only for the ATE estimate itself but also for the variance estimate that is used to construct the CIs. In other words, both components must be computed in a differentially private manner to ensure privacy.
@@
So far, there are only few works on treatment effect estimation under DP [e.g., 29, 44, 52]. Existing works mainly focus on estimating causal quantities as point estimates, whereas privacy-preserving CIs are missing. As discussed above, extending existing methods for ATE estimation under DP to construct differentially private CIs is not readily feasible. Other studies focus on privacy-preserving CIs [e.g., 9, 24], yet outside of treatment effect estimation and, therefore, cannot be directly applied to causal quantities such as the ATE. Guha and Reiter [29] and Ohnishi and Awan [52] present private CIs for the ATE, but in different settings (e.g., by making restrictive assumptions on the data-generating process, by being limited to linear models, and/or by focusing on a different definition of DP). To our knowledge, a method for computing valid CIs for the ATE under ($\varepsilon,\delta$)-DP is missing.
@@
Here, we develop PrivATE, the first framework designed to compute valid CIs for the ATE under ($\varepsilon,\delta$)-DP (see Figure 1). Our framework is built upon the augmented inverse probability weighting (AIPW) estimator, which offers several practical advantages such as being model-agnostic, doubly robust, and Neyman-orthogonal [54, 64]. Of note, model-agnostic means that our framework can be used together with arbitrary machine learning models such as neural networks. To obtain the CIs, we further introduce a novel variance estimator for the ATE that is based on truncated output perturbation and satisfies DP. This allows us to construct valid, model-agnostic CIs even in the presence of privatization noise.
@@ 2 Related work
Differentially private ATE estimation: Research on privacy mechanisms for ATE estimation is scarce. Lee et al. [44] design a privacy-preserving Horvitz-Thompson type estimator by privatizing the propensity score function. However, due to the employed regularization, the privatized ATE estimate is eventually biased, which prevents the construction of valid CIs. Guha and Reiter [29] provide a DP method for the estimation of ATEs of binary outcomes. Other works have analyzed private treatment effect estimation in other DP regimes like local DP [2, 51] or label DP [37], or focus on other causal quantities [e.g., 52, 60, 67]. However, as explained in our introduction, computing differentially private CIs for the ATE is non-trivial.
Differentially private CIs for non-causal quantities: Several works focus on the construction of differentially private CIs, yet outside of causal inference. Here, one stream constructs CIs for quantities estimated through parametric models where the underlying population is known and/or normally distributed [e.g., 40, 19, 24]. Other works have focused on constructing CIs for the parameters of a trained model, e.g., linear regression [9, 61]. A more general method has been proposed to construct asymptotic confidence intervals for parameters of any model trained by empirical risk minimization [65]; however, this method imposes strong assumptions on the convexity and the smoothness of the loss function, making it incompatible with many machine learning models. In sum, the above works focus on non-causal quantities and, therefore, are not directly applicable to ATE estimation. Therefore, we must first derive an asymptotically normal differentially private ATE estimate before we can then obtain private CIs through our framework.
@@
CIs for the ATE: Several works construct CIs for the ATE based on $\sqrt{n}$-consistent estimators with asymptotic coverage guarantees [8, 34]. Some works also focus on finite-sample CIs. However, these require significantly stronger assumptions on the data [4]. Finally, [66] uses prediction-powered inference to derive CIs for the ATE from multiple datasets. However, no method ensures DP.
Literature on privacy mechanisms for CIs of the ATE is scarce. We are aware of only two works [29, 52], yet both works have clear limitations. Guha and Reiter [29] focus on the WATE (i.e., the weighted ATE, which, in principle, can be transformed into the ATE), but the method puts restrictive assumptions on the data-generating process, because of which the method is later not applicable. [52] is limited to a sieve logistic regression and is not designed to deal with neural networks.
Importantly, both works focus on a different privacy definition, namely, pure $\varepsilon$-DP, while we focus on the more common ($\varepsilon,\delta$)-DP. On top of that, unlike our PrivATE framework, none of the two methods benefit from the favorable properties of being doubly robust and Neyman-orthogonal.
Research gap: To the best of our knowledge, we are the first to provide a general machine learning framework to construct valid CIs for the ATE under ($\varepsilon,\delta$)-DP.
@@ 3 Problem formulation
Notation: We denote random variables by capital letters $X$ and their realizations by small letters $x$. We refer to the probability distribution over $X$ by $P_{X}$, where we omit the subscript whenever it is apparent from the context. The probability mass function for discrete $X$ is given by $P(x)=P(X=x)$ and the probability density function w.r.t. the Lebesgue measure by $p(x)$.
Setting: We consider a dataset $D:=\{(X_{i},A_{i},Y_{i})\}_{i=1,\ldots,n}$, consisting of observed confounders $X$, a binary treatment $A\in\{0,1\}$, and a bounded outcome $Y\in\mathcal{Y}$, where $Z_{i}:=(X_{i},A_{i},Y_{i})\sim P$ i.i.d., $Z_{i}\in\mathcal{Z}$, and $\mathcal{X},\mathcal{Y}$ have bounded domains. We use the potential outcomes framework [57] and denote the potential outcome of intervention $a$ by $Y(a)$. Furthermore, we define the propensity score as $\pi(x):=P(A=1\mid X=x)$ and the outcome function as $\mu(x,a):=\mathbb{E}[Y\mid X=x,A=a]$.
Objective: We estimate the average treatment effect (ATE) $\tau$ under ($\varepsilon,\delta$)-DP and construct CIs for $\tau:=\mathbb{E}[Y(1)-Y(0)]$ so that the CIs (i) capture both estimation and privatization uncertainty and (ii) retain the DP guarantees. We make the standard assumptions in causal inference: positivity, consistency, and unconfoundedness [e.g., 17, 57]. Then, the ATE is identified as $\tau=\mathbb{E}[Y\mid A=1]-\mathbb{E}[Y\mid A=0]$. We provide an introduction to ATE estimation and an overview of the corresponding estimators in Supplement A.
In the following, we derive differentially private CIs for the ATE estimated via the AIPW estimator. The AIPW estimator has several advantages [55, 64]: it is unbiased and asymptotically normally distributed. Furthermore, the AIPW estimator is doubly robust and Neyman-orthogonal, making it insensitive to small errors in the estimated nuisances $\hat{\mu}$ and $\hat{\pi}$. However, as we show in the following, it is non-trivial to construct private CIs from the AIPW estimator.
@@ 3.1 Differential privacy
To ensure the privacy of sensitive information, ($\varepsilon,\delta$)-differential privacy (DP) guarantees that the inclusion or exclusion of any individual data point does not affect the estimated outcome by more than a pre-specified threshold defined by the privacy budget $\varepsilon$ and $\delta$ [20, 21]. Specifically, the probability density of any outcome $y$ on dataset $D\in\mathcal{Z}^{n}$ must be $\varepsilon$-indistinguishable from the probability density of the same outcome $y$ on a neighboring dataset $D^{{}^{\prime}}\in\mathcal{Z}^{n}$ with a probability of at least $1-\delta$.
Definition 3.1 (Differential privacy [21]).
Two datasets $D$ and $D^{{}^{\prime}}$ are called neighbors, denoted as $D\sim D^{{}^{\prime}}$, if their Hamming distance equals one, i.e., $d_{\mathrm{H}}(D,D^{{}^{\prime}})=1$. A function $\mathbf{f}_{D}:\mathcal{X}^{d}\mapsto\mathbb{R}^{d}$ trained on dataset $D$ is $(\varepsilon,\delta)$-differentially private if, for all neighboring datasets $D$, $D^{{}^{\prime}}\in\mathcal{Z}^{n}$, all measurable $S\subseteq\mathbb{R}^{d}$ and all $\mathbf{x}\in\mathcal{X}^{d}$, it holds that
$P(\mathbf{f}_{D}(\mathbf{x})\in S)\leq\exp(\varepsilon)\cdot P(\mathbf{f}_{D^{{}^{\prime}}}(\mathbf{x})\in S)+\delta.$ (1)
DP can be achieved by employing the so-called Gaussian privacy mechanisms [22], where the aim is to perturb the prediction by adding appropriately calibrated zero-centered Gaussian noise so that two predictions resulting from neighboring databases cannot be differentiated [e.g., 15, 68].
Definition 3.2 (Gaussian privacy mechanism [22]).
Let $\mathbf{f}:\mathcal{X}^{d}\mapsto\mathbb{R}^{d}$ and let $\Delta_{2}(\mathbf{f})=\sup_{D\sim D^{{}^{\prime}}\!\!,\,\mathbf{x}\in\mathcal{X}^{d}}||\mathbf{f}_{D}(\mathbf{x})-\mathbf{f}_{D^{{}^{\prime}}}(\mathbf{x})||$ denote the $l_{2}$-sensitivity if $\mathbf{f}$. Let $\mathbf{U}\sim\mathcal{N}(0,\sigma^{2})$ for $\sigma\geq\frac{1}{\varepsilon}\sqrt{2\ln{(1.25/\delta)}}\cdot\Delta_{2}(\mathbf{f})$. Then, $\mathbf{f}^{\textrm{{DP}}}_{D}(\mathbf{x})=\mathbf{f}_{D}(\mathbf{x})+\mathbf{U}$ is $(\varepsilon,\delta)$-differentially private.
@@ 3.2 Problem statement
Uncertainty quantification for the ATE is commonly presented in terms of confidence intervals. [^3] For an asymptotically consistent ATE estimate $\hat{\tau}$ and its estimated variance $\hat{\sigma}^{2}$, the CI for a confidence level $(1-\alpha)$, $\alpha\in(0,1)$, is given by $\mathrm{CI}=\Big{[}\hat{\tau}\pm\Phi^{-1}\Big{(}1-\frac{\alpha}{2}\Big{)}\sqrt{\frac{\hat{\sigma}^{2}}{n}}\Big{]},$ where $\Phi^{-1}\Big{(}1-\frac{\alpha}{2}\Big{)}$ denotes the critical value given by the respective quantile of the standard normal distribution. In our work, we aim at differentially private uncertainty quantification of the ATE. Specifically, we aim to derive a $(\varepsilon,\delta)$-differentially private estimate $\hat{\tau}_{\textrm{{DP}}}$ and the corresponding differentially private CI
@@ 4.2 Mathematical preliminaries
Our work exploits a connection between model robustness and privacy in terms of the influence function of the underlying estimator following former work in this field [e.g., 21, 60].
@@
The IF of the AIPW learner is given by the influence function of the final mean estimation in the second step. Specifically, the influence stemming from training the nuisance estimators $\pi$ and $\mu$ can be neglected during the privatization step if the estimators are sufficiently stable. [^4] We require the base estimators for $\mu$ and $\pi$ to be sufficiently stable in terms of uniform stability [11]. This is necessary to provide fast convergence of the (possibly misspecified) nuisance estimators. It is worth noting that many common ML methods fulfill this criterion.
@@
Remark: To avoid stability guarantees, one can alternatively privatize the nuisances by a DP-method of choice. The privatization neither affects the efficiency of the estimator [60] nor the validity of the CIs. Specifically, we split the dataset $D$ into two disjoint subsets $D_{1}$ and $D_{2}$. On $D_{1}$, we estimate $\hat{\pi},\hat{\mu}$ under $(\varepsilon_{1}/2,\delta_{1}/2)$-DP, which are then employed to estimate $\hat{\tau}$ on $D_{2}$. Note that data splitting is common to prevent overfitting and guarantee efficiency of the estimators [e.g., 26, 64].

### v2

@@ Abstract
The average treatment effect (ATE) is widely used to evaluate the effectiveness of drugs and other medical interventions. In safety-critical applications like medicine, reliable inferences about the ATE typically require valid uncertainty quantification, such as through confidence intervals (CIs). However, estimating treatment effects in these settings often involves sensitive data that must be kept private. In this work, we present PrivATE, a novel machine learning framework for computing CIs for the ATE under differential privacy. Specifically, we focus on deriving valid privacy-preserving CIs for the ATE from observational data. Our PrivATE framework consists of three steps: (i) estimating the differentially private ATE through output perturbation; (ii) estimating the differentially private variance in a doubly robust manner; and (iii) constructing the CIs while accounting for the uncertainty from both the estimation and privatization steps. Our PrivATE framework is model agnostic, doubly robust, and ensures valid CIs. We demonstrate the effectiveness of our framework using synthetic and real-world medical datasets. To the best of our knowledge, we are the first to derive a general, doubly robust framework for valid CIs of the ATE under ($\varepsilon,\delta$)-differential privacy.
@@ 1 Introduction
Estimating the average treatment effect (ATE) from observational data is highly relevant for evaluating the effectiveness of drugs and other medical interventions (e.g., Ballmann, 2015; Buell et al., 2024; Feuerriegel et al., 2024). To ensure the reliability of ATE estimates, one often needs to “move beyond the mean” by accounting for the estimation uncertainty (Heckman et al., 1997; Kneib et al., 2023). This is typically achieved through confidence intervals (CIs), which provide a range within which the true effect is likely to lie (Neyman, 1937). Medical studies frequently report 95% CIs to show the precision of the estimated treatment effects. For example, the clinical trial for the Moderna COVID-19 vaccine specified that the effectiveness must meet a predefined threshold based on the 95% CI, and, ultimately, the trial reported a 95% CI of 89.3% to 96.8% (Baden et al., 2021).
Estimating ATEs in medicine typically involves sensitive patient data (Brothers and Rothstein, 2015), which must be kept private. Several regulations have been introduced to enforce data privacy, including the US Health Insurance Portability and Accountability Act (HIPAA) and the EU General Data Protection Regulation (GDPR). One approach to analyse sensitive data is to use privacy mechanisms such as differential privacy (DP), which enables population-level inference while protecting individual information (Dwork, 2006). Due to strong theoretical guarantees, especially ($\varepsilon,\delta$)-DP has received increasing attention (e.g., Abadi et al., 2016; Bassily et al., 2014).
In this work, our aim is to compute differentially private confidence intervals for the ATE. However, this is non-trivial: constructing CIs solely based on the private ATE estimate does not suffice to ensure DP of the entire CI, as estimating the variance required for constructing CIs involves additional queries to the underlying data. Thus, it is necessary to develop privacy-preserving mechanisms not only for the ATE estimate itself but also for the variance estimate. In other words, both components must be computed in a differentially private manner to ensure privacy.
@@
So far, there are only a few works on treatment effect estimation under DP (e.g., Guha and Reiter, 2024; Lee et al., 2019; Ohnishi and Awan, 2024). Existing works mainly focus on estimating causal quantities as point estimates, whereas privacy-preserving CIs are missing. As discussed above, extending existing methods for ATE estimation under DP to construct differentially private CIs is not readily feasible. Other studies focus on privacy-preserving CIs (e.g., Barrientos et al., 2019; Ferrando et al., 2022), yet outside of treatment effect estimation and, therefore, cannot be directly applied to causal quantities such as the ATE. Guha and Reiter (2024) and Ohnishi and Awan (2024) present private CIs for the ATE, but in different settings (e.g., by making restrictive assumptions on the data-generating process, by being limited to linear models, and/or by focusing on a different definition of DP). To the best of our knowledge, a method for computing valid CIs for the ATE under ($\varepsilon,\delta$)-DP is missing.
@@
Here, we develop PrivATE, the first framework designed to compute valid CIs for the ATE under ($\varepsilon,\delta$)-DP (see Figure 1). Our framework is built upon the augmented inverse probability weighting (AIPW) estimator, which offers several practical advantages such as being model-agnostic and doubly robust (Robins and Rotnitzky, 1995; Wager, 2024). Notably, model-agnostic means that our framework can be used in conjunction with any machine learning model. To obtain the CIs, we further introduce a novel variance estimator for the ATE satisfying DP. This allows us to construct valid, model-agnostic CIs even in the presence of privatization noise.
@@ 2 Related work
Differentially private ATE estimation: Research on privacy mechanisms for ATE estimation is scarce. Lee et al. (2019) design a privacy-preserving Horvitz-Thompson type estimator by privatizing the propensity score function. However, due to the employed regularization, the privatized ATE estimate is eventually biased, which prevents the construction of valid CIs. Guha and Reiter (2024) provide a DP method for the estimation of ATEs of binary outcomes. Other works have analyzed private treatment effect estimation in other DP regimes like local DP (Agarwal and Singh, 2024; Ohnishi and Awan, 2023) or label DP (Javanmard et al., 2024), or focus on other causal quantities (e.g., Ohnishi and Awan, 2024; Schröder et al., 2025; Yao et al., 2024). However, as explained in our introduction, computing differentially private CIs for the ATE is non-trivial.
Differentially private CIs for non-causal quantities: Several works focus on the construction of differentially private CIs, yet outside of causal inference. Here, one stream constructs CIs for quantities estimated through parametric models where the underlying population is known and/or normally distributed (e.g., Karwa and Vadhan, 2018; D’Orazio and Honaker, 2015; Ferrando et al., 2022). Other works have focused on constructing CIs for the parameters of a trained model, e.g., linear regression (Barrientos et al., 2019; Sheffet, 2017). A more general method has been proposed to construct asymptotic confidence intervals for parameters of any model trained by empirical risk minimization (Wang et al., 2019); however, this method imposes strong assumptions on the convexity and the smoothness of the loss function, making it incompatible with many machine learning models. In sum, the above works focus on non-causal quantities and, therefore, are not directly applicable to ATE estimation. Therefore, we must first derive an asymptotically normal differentially private ATE estimate before we can then obtain private CIs through our framework.
@@
CIs for the ATE: Several works construct CIs for the ATE based on $\sqrt{n}$-consistent estimators with asymptotic coverage guarantees (Bang and Robins, 2005; Hirano et al., 2003). Some works also focus on finite-sample CIs. However, these require significantly stronger assumptions on the data (Aronow et al., 2021). Finally, Wang et al. (2025) uses prediction-powered inference to derive CIs for the ATE from multiple datasets. However, no method ensures DP.
Literature on privacy mechanisms for CIs of the ATE is scarce. We are aware of only two works (Guha and Reiter, 2024; Ohnishi and Awan, 2024), yet both works have clear limitations. Guha and Reiter (2024) focus on the WATE (i.e., the weighted ATE, which, in principle, can be transformed into the ATE), but the method puts restrictive assumptions on the data-generating process, because of which the method is later not applicable. The approach by Ohnishi and Awan (2024) is limited to a sieve logistic regression and is not designed to deal with arbitrary ML models. Importantly, both works focus on a different privacy definition, namely, pure $\varepsilon$-DP, while we focus on the more common ($\varepsilon,\delta$)-DP.
Research gap: To our knowledge, we are the first to provide a general machine learning framework to construct valid CIs for the ATE under ($\varepsilon,\delta$)-DP.
@@ 3 Problem formulation
Notation: We denote random variables by capital letters $X$ and their realizations by small letters $x$. We refer to the probability distribution over $X$ by $P_{X}$, where we omit the subscript if it is apparent from the context. The probability mass function for discrete $X$ is given by $P(x)=P(X=x)$ and the probability density function w.r.t. the Lebesgue measure by $p(x)$.
Setting: We consider a dataset $D:=\{(X_{i},A_{i},Y_{i})\}_{i=1,\ldots,n}$, consisting of observed confounders $X$, a binary treatment $A\in\{0,1\}$, and a bounded outcome $Y\in\mathcal{Y}$, where $Z_{i}:=(X_{i},A_{i},Y_{i})\sim P$ i.i.d., $Z_{i}\in\mathcal{Z}$, and $\mathcal{X},\mathcal{Y}$ have bounded domains. We use the potential outcomes framework (Rubin, 2005) and denote the potential outcome of intervention $a$ by $Y(a)$. Furthermore, we define the propensity score as $\pi(x):=P(A=1\mid X=x)$ and the outcome function as $\mu(x,a):=\mathbb{E}[Y\mid X=x,A=a]$.
Objective: We estimate the average treatment effect (ATE) $\tau$ under ($\varepsilon,\delta$)-DP and construct CIs for $\tau:=\mathbb{E}[Y(1)-Y(0)]$ so that the CIs (i) capture both estimation and privatization uncertainty and (ii) retain the DP guarantees. We make the standard assumptions in causal inference: positivity, consistency, and unconfoundedness (e.g., Curth and van der Schaar, 2021; Rubin, 2005). Then, the ATE is identified as $\tau=\mathbb{E}[Y\mid A=1]-\mathbb{E}[Y\mid A=0]$. We provide an introduction to ATE estimation and an overview of the corresponding estimators in Supplement A.
In the following, we derive differentially private CIs for the ATE estimated via the AIPW estimator. The AIPW estimator has several advantages (Robins et al., 1994; Wager, 2024): it is unbiased and asymptotically normally distributed if at least one of the nuisances is correctly specified. Furthermore, the AIPW estimator is doubly robust and Neyman-orthogonal, making it insensitive to small errors in the estimated nuisances $\hat{\mu}$ and $\hat{\pi}$ for nuisances converging fast enough to the oracle, i.e., that convergence rates are in $O(n^{-\beta_{\mu}})$ and $O(n^{-\beta_{\pi}})$, respectively, with $\beta_{\mu}+\beta_{\pi}\geq\frac{1}{2}$. We make this standard assumption of sufficiently fast convergence throughout our work (e.g., Wager, 2024; Wang et al., 2025).
@@ 3.1 Differential privacy
To ensure the privacy of sensitive information, ($\varepsilon,\delta$)-differential privacy (DP) guarantees that the inclusion or exclusion of any individual data point does not affect the estimated outcome by more than a pre-specified threshold defined by the privacy budget $\varepsilon$ and $\delta$ (Dwork, 2006; Dwork and Lei, 2009). Specifically, the probability density of any outcome $y$ on dataset $D\in\mathcal{Z}^{n}$ must be $\varepsilon$-indistinguishable from the probability density of the same outcome $y$ on a neighboring dataset $D^{{}^{\prime}}\in\mathcal{Z}^{n}$ with a probability of at least $1-\delta$.
Definition 3.1 (Differential privacy (Dwork and Lei, 2009)).
Two datasets $D$ and $D^{{}^{\prime}}$ are called neighbors, denoted as $D\sim D^{{}^{\prime}}$, if their Hamming distance equals one, i.e., $d_{\mathrm{H}}(D,D^{{}^{\prime}})=1$. A function $\mathbf{f}_{D}:D\mapsto\mathbb{R}^{d}$ trained on dataset $D$ is $(\varepsilon,\delta)$-differentially private if, for all neighboring datasets $D$, $D^{{}^{\prime}}\in\mathcal{Z}^{n}$ and all measurable $S\subseteq\mathbb{R}^{d}$, it holds that
$P(\mathbf{f}_{D}\in S)\leq\exp(\varepsilon)\cdot P(\mathbf{f}_{D^{{}^{\prime}}}\in S)+\delta.$ (1)
DP can be achieved by employing the so-called Gaussian privacy mechanisms (Dwork and Roth, 2014), where the aim is to perturb the prediction by adding appropriately calibrated zero-centered Gaussian noise so that two predictions resulting from neighboring databases cannot be differentiated (e.g., Chaudhuri et al., 2011; Zhang et al., 2022).
Definition 3.2 (Gaussian privacy mechanism (Dwork and Roth, 2014)).
[…]
@@ 3.2 Problem statement
Uncertainty quantification for the ATE is commonly presented in terms of confidence intervals. [^3] For an asymptotically consistent ATE estimate $\hat{\tau}$ and its estimated variance $\hat{\sigma}^{2}$, the CI for a confidence level $(1-\alpha)$, $\alpha\in(0,1)$, is given by $\mathrm{CI}=\Big[\hat{\tau}\pm\Phi^{-1}\Big(1-\frac{\alpha}{2}\Big)\sqrt{\frac{\hat{\sigma}^{2}}{n}}\Big],$ where $\Phi^{-1}\Big(1-\frac{\alpha}{2}\Big)$ denotes the critical value given by the respective quantile of the standard normal distribution. In our work, we aim at differentially private uncertainty quantification of the ATE. Specifically, we aim to derive an $(\varepsilon,\delta)$-differentially private estimate $\hat{\tau}_{\textrm{{DP}}}$ and the corresponding differentially private CI
@@ 4.2 Mathematical preliminaries
$\hat{\tau}=\frac{1}{n}\sum_{i=1}^{n}\Bigg[\hat{\mu}(X,1)-\hat{\mu}(X,0)+\frac{Y_{i}-\hat{\mu}(X,1)}{\hat{\pi}(X_{i})}A_{i}$ (3)
$-\frac{Y_{i}-\hat{\mu}(X,0)}{1-\hat{\pi}(X_{i})}(1-A_{i})\Bigg].$ (4)
Our work exploits a connection between model robustness and privacy in terms of the influence function of the underlying estimator following former work in this field (e.g., Dwork and Lei, 2009; Schröder et al., 2025).
@@
We make the following assumptions on the nuisance functions: The nuisance functions are (i) bounded, (ii) estimated at rates $n^{-\beta_{\mu}}$ and $n^{-\beta_{\pi}}$ with $\beta_{\mu}+\beta_{\pi}\geq\frac{1}{2}$, and (iii) in a local neighborhood of the true nuisance functions, i.e., there exists $\lambda_{n}$ decreasing in the sample size $n$, s.t. $\rVert\hat{\eta}-\eta_{0}\lVert{\infty}\leq\lambda_{n}$. Then, the IF of the AIPW learner is dominated by the influence function of the final mean estimation in the second step. The influence stemming from training the nuisance estimators $\hat{\pi}$ and $\hat{\mu}$ is negligible in the privatization step.
@@
Remark: Assumptions (i) states a weak regularity condition, restricting the nuisance predictions to fall into the bounded domains $\mathcal{A},\mathcal{Y}$. Assumption (ii) postulates the standard convergence rate requirement for Neyman-orthogonality introduced in Section 3. Finally, assumption (iii) requires that the nuisances are estimated sufficiently well. Together with the Neyman-orthogonality, it renders the influence stemming from the first stage estimation negligible for privatization.
@@
Remark: To avoid making the local neighborhood assumption, one can alternatively privatize the nuisances by a DP method of choice. The privatization neither affects the efficiency of the estimator (Schröder et al., 2025) nor the validity of the CIs. Specifically, we split the dataset $D$ into disjoint subsets $D_{1}$ and $D_{2}$. On $D_{1}$, we estimate $\hat{\pi},\hat{\mu}$ under $(\varepsilon_{1}/2,\delta_{1}/2)$-DP, which are then employed to estimate $\hat{\tau}$ on $D_{2}$. Note that data splitting is common to prevent overfitting of the estimators (e.g., Foster and Syrgkanis, 2019; Wager, 2024).